# Boundary vertex algebras for 3d $\mathcal{N}=4$ rank-0 SCFTs

### v2

@@ 1 Introduction
The underlying intuition is that the $HT$ twist already captures IR information, and should therefore be able to see the emergent supersymmetry, cf. GRW . The existence of extra supercurrents extending ${\cal N}=2$ to ${\cal N}=4$, and therefore of two topological twists in the IR, manifests itself in the $HT$ twist as the presence of two distinguished operators whose descendants can be used to deform the holomorphic-topological theory to two theories that are fully topological. Roughly speaking, this deformation can be thought of as turning on a suitable superpotential. From a BV/BRST perspective, this deforms the differential $Q$ of the $HT$-twisted theory, which is the sum of the BV/BRST differential and the twisting supercharge $Q=Q_{BRST}+Q_{HT}$, to $Q_{A/B}=Q+\delta_{A/B}$. The effect of such a superpotential deformation on the boundary vertex algebra can be understood using the analysis of CDG . As we review below, see also twistedN=4 ; GRW , the second operator plays a distinguished role after deforming: it trivializes (makes homotopically trivial) the bulk’s dependence on the holomorphic coordinate and oftentimes realizes an action of holomorphic vector fields at the boundary, which is expected to extend to an action of the Virasoro algebra (thus producing a VOA). Importantly, the form of the superpotential used to deform to topological theories instructs which boundary conditions are deformable in the sense of CG ; BLS .
@@
where we have $r$ copies of a $U(1)$ gauge theory with a chiral multiplet of charge $1$, a suitable matrix $K_{r}$ of mixed Chern-Simons level, and $V_{{\mathfrak{m}}_{i}}$ are a basis of gauge-invariant half-BPS monopole operators. In the IR, a $U(1)\subset(U(1)^{r})_{\text{top}}$ subgroup of the topological symmetry mixes with the $\mathcal{N}=2$ R-symmetry $U(1)_{R}$ to form the $\mathcal{N}=4$ R-symmetry group $(SU(2)_{H}\times SU(2)_{C})/\mathbb{Z}_{2}$; the remaining topological symmetries are broken by the monopoles operators, in agreement with the rank-0 expectation. One of the results of GKS was that certain half indices of these theories, in the limits appropriately reproducing the $A$ and $B$ twists, appear to be vector-valued modular forms closely related to characters of the Virasoro minimal models $M(2,2r+3)$. As we now explain, our main motivation to study these boundary VOAs came from the observation of these modular phenomena as well as the appearance of Virasoro minimal models. We note that these minimal models also arise in the context of 4d Argyres-Douglas theories Beem:2017ooy and their reduction to 3d Dedushenko:2018bpp ; MD:ToAppear .
@@
    Based on examples Costello:2018swh , the Higgs branch conjecture of Beem-Rastelli in 4d, as well as free-field realizations in 3d by Beem and the first author Beem:2023dub , the Higgs/Coulomb branch chiral ring is expected to be encoded into (the reduced part of) Zhu’s $C_{2}$ algebra $R_{\cal V}$ of the boundary VOA ${\cal V}$; in Beem:2023dub it was conjectured that for a large class of A-twisted abelian gauge theories, the associated variety $\mathrm{Specm}(R_{\cal V})$ is indeed the Higgs branch.
@@
Thus, based on these expectations, it is natural to strengthen a bit the rank-0 conditions, and study the case where 1) $R_{\cal V}$ is finite-dimensional and 2) there are no non-trivial self-extensions of the vacuum module. The first condition already has important consequences. In fact, it is known that under some technical assumptions (that the VOA is finitely strongly generated and non-negatively graded, which usually holds) the $C_{2}$-cofiniteness condition (and as a consequence, the zero-dimensionality of the associated variety) implies that the vertex algebra is lisse arakawa2012remark . This further implies that its characters indeed enjoy modular properties, see e.g. Zhu . Therefore, taking seriously the first expectation means that rank-0 theories have a good chance of admitting lisse boundary VOAs, and in turn, of having half-indices that enjoy modular properties.
@@
We consider the higher-rank theories ${\cal T}_{r}$ in Section 5, where a direct description of the boundary VOA becomes somewhat more difficult. We argue that there is a Dirichlet boundary condition Dir${}^{(r)}$ that is compatible with the $B$-twist and provide evidence for the algebra of boundary local operators being identified with the simple affine VOA $L_{r}(\mathfrak{osp}(1|2))$. We finish Section 5 by comparing properties of the category of $L_{r}(\mathfrak{osp}(1|2))$ modules with the analysis of GKS , finding compatible results. The form of our half-indices can be viewed as fermionic sum representations of $L_{r}(\mathfrak{osp}(1|2))$ characters, which may be of independent interest. Finally, in Appendix A we make use of a mechanism introduced in Section 4 to propose a twisted 3d ${\cal N}=2$ TQFT and boundary condition thereof that realizes the minimal model $M(3,4)$.
@@ 3.1 The $HT$ twist of an ${\cal N}=4$ theory
We start by presenting some general features of the $HT$ twist of an ${\cal N}=4$ theory, following the conventions of twistedN=4 . The 3d ${\cal N}=4$ supersymmetry algebra has $8$ real spinor supercharges $Q^{a\dot{a}}_{\alpha}$, where $a,\dot{a}$ are $SO(4)_{R}\simeq(SU(2)_{H}\times SU(2)_{C})/\mathbb{Z}_{2}$ $R$-symmetry spinor indices and $\alpha$ is a $SU(2)_{E}$ Lorentz spinor index, and the following anti-commutators
@@ 3.2 Supercurrents and superpotentials
We now see how to deform the $Q_{HT}$ twist of an ${\cal N}=4$ theory to the topological $Q_{A/B}$ twist. First, we redefine the twisted spin and the $R$-charge to be those of the topological $A/B$ twist. This ensures the operator $\Theta_{A/B}$ has the necessary quantum numbers; for example, in the $A$-twist we take the $R$-charge generated by $R-S=2R_{C}$ and the twisted spin generated by $J+\tfrac{1}{2}S=R_{H}-J_{3}$. This change in twisted spin is accounted for in a modification of the higher stress tensor $G\to G_{A/B}=G\pm\tfrac{1}{2}\partial_{z}{\cal S}$.
@@ 4 A boundary vertex algebra for ${\cal T}_{\text{min}}$
are the operators realizing the deformations to the $A$ and $B$ twist; the charge for the $U(1)_{S}$ symmetry is identified with (the negative of the) monopole number. In view of the analysis in GKS , there is a natural candidate deformable boundary condition: we expect that the Dirichlet boundary condition $\text{Dir}=({\cal D},D)$ imposing $(0,2)$ Dirichlet boundary conditions on both the Chern-Simons vector multiplet and the chiral multiplet is deformable to the topological $B$-twist. Indeed, Dirichlet boundary conditions for the chiral multiplet impose $\phi|=0$ and leave $\overline{\lambda}_{-}|$ unconstrained, therefore $\Theta_{B}|=0$. Moreover, the Dirichlet boundary conditions on the vector multiplet further imply the bulk local operator $\Theta_{A}$ is unconstrained on Dir. The local operators on Dir are counted by the half-index GGP ; DGP ; Yoshida:2014ssa , which is defined by
@@
Based on the discussion in Section 5, the generic Dirichlet boundary condition considered in GKS is deformable for the $A$ twist of ${\cal T}_{\text{min}}$ – although $\Theta_{A}$ survives on Dir, for generic Dirichlet boundary conditions there is an additional differential that makes $\Theta_{A}$ trivial in cohomology. Moreover, the operator $\Theta_{B}$ is non-vanishing on this generic Dirichlet boundary condition, whereas it vanishes on Dir. We leave a deeper treatment of the algebra of local operators on the generic Dirichlet boundary condition of GKS for future work. [^3] Another candidate for a deformable boundary condition for the $A$ twist is $({\cal D},N)$, imposing $(0,2)$ Dirichlet boundary conditions on the vector multiplet and $(0,2)$ Neumann boundary conditions for the chiral multiplet. The Neumann boundary conditions on the chiral multiplet impose $\overline{\lambda}_{-}|=0$ which implies $\Theta_{A}=0$. However, the boundary theory has anomalous $U(1)_{H}$, which is incompatible with the $A$-twist. This can be remedied by adding a boundary fermion with $R$-charge $1$ and $U(1)_{\partial}$ charge $-1$. The $HT$-twisted half-index for $({\cal D},N)$ takes the form $I\!\!I_{({\cal D},N)}(q;y,\eta,\nu)$ $=\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{q^{\scriptsize{\frac{1}{2}}\mathfrak{m}^{2}}y^{\mathfrak{m}}[(-q^{\scriptsize{\frac{1}{2}}})^{\nu}\eta]^{\mathfrak{m}}}{(q;q)_{\infty}(yq^{\mathfrak{m}};q)_{\infty}}$ $=\bigg{(}\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{(-1)^{\mathfrak{m}}q^{\scriptsize{\frac{1}{2}}\mathfrak{m}(\mathfrak{m}+\nu)}y^{\mathfrak{m}}\eta^{\mathfrak{m}}}{(q;q)_{\infty}}\bigg{)}\bigg{(}\sum_{n\geq 0}\frac{(-1)^{n}q^{-\scriptsize{\frac{1}{2}}n(n+\nu)}\eta^{-n}}{(q;q)_{n}}\bigg{)}\ .$ Deforming to the $A$-twist amounts to sending $\eta\to 1$ and $\nu\to-1$, leading to $I\!\!I_{A}(q;y)=I\!\!I_{({\cal D},N)}(q;y,1,-1)=\bigg{(}\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{(-1)^{\mathfrak{m}}q^{\scriptsize{\frac{1}{2}}\mathfrak{m}(\mathfrak{m}-1)}y^{\mathfrak{m}}}{(q;q)_{\infty}}\bigg{)}\bigg{(}\sum_{n\geq 0}\frac{(-1)^{n}q^{\scriptsize{\frac{1}{2}}n(1-n)}}{(q;q)_{n}}\bigg{)}\ ,$ which has unbounded powers of $q$. If we add a boundary Fermi multiplet with $U(1)_{\partial}$ and $U(1)_{R}$ charges $-1$ and $1$ respectively, as required to cancel the boundary anomaly, we need to multiply $\mathfrak{m}$th summand in the half-index by $(yq^{\mathfrak{m}};q)_{\infty}(q^{1-\mathfrak{m}}y^{-1};q)_{\infty}~{}.$ This has the pleasing effect of making the $q$-expansion bounded from below.
@@ 4.1.1 The $HT$-twisted boundary vertex algebra
The level of the abelian current reflects the effective Chern-Simons $k_{eff}=2$. Each of these operators has spin $1$.
@@
$\sigma_{\mathfrak{m}}(B_{n})=B_{n}-2\mathfrak{m}\delta_{n+\mathfrak{m},0}\ ,\qquad\sigma_{\mathfrak{m}}(\overline{\lambda}_{n})=\overline{\lambda}_{n+\mathfrak{m}}\ ,$ (45)
@@
We note that the subalgebra generated by $B$ and the operators corresponding to spectral flows of the vacuum $\sigma_{\mathfrak{m}}(|0\rangle)$ can be identified with a rank-1 lattice VOA at level $2$, i.e. based on the lattice $\sqrt{2}\mathbb{Z}$. This particularly important lattice VOA can be identified with (the simple quotient of) an affine $\mathfrak{sl}(2)$ current algebra at level $1$. In particular, if we identify $B\leftrightarrow h$, $e\leftrightarrow\sigma_{1}(|0\rangle)$, and $f\leftrightarrow\sigma_{-1}(|0\rangle)$, we get the following non-singular OPEs:
@@
From the above spectral flow analysis we also conclude that the OPE of $f$ and $\overline{\lambda}$ is non-singular, whence it is a lowest weight vector of a doublet for this affine $\mathfrak{sl}(2)$; the state $\overline{\lambda}_{0}\sigma_{1}(|0\rangle)=\sigma_{1}(\overline{\lambda}_{-1}|0\rangle)$ fills out the remaining states in this doublet and we identify it with the restriction of the bare monopole $V_{1}|$. (The bare monopole $V_{-1}|$ is identified with $f$.) If we denote the operator dual to this state $\theta_{+}$ and rename $\overline{\lambda}=\theta_{-}$, we arrive at the following non-singular OPE
@@
Associativity then constrains the OPEs with the remaining $\mathfrak{sl}(2)$ currents to be
$h(z)\theta_{\pm}(w)\sim\frac{\pm\theta_{\pm}(w)}{z-w}\qquad f(z)\theta_{+}(w)\sim\frac{-\theta_{-}(w)}{z-w}$ (50)
Finally, we note that the OPE of $\theta_{+}$ and $\theta_{-}$ is non-singular; this follows from the action of $\overline{\lambda}_{n}$ on $\sigma_{1}(\overline{\lambda}_{-1}|0\rangle)$. We expect that the algebra of boundary local operators is (strongly, but not freely) generated by the bosonic operators $e,f,h$ and the fermionic operators $\theta_{\pm}$.

### v3

@@ 1 Introduction
The underlying intuition is that the $HT$ twist already captures IR information, and should therefore be able to see the emergent supersymmetry, cf. GRW . The existence of extra supercurrents extending ${\cal N}=2$ to ${\cal N}=4$, and therefore of two topological twists in the IR, manifests itself in the $HT$ twist as the presence of two distinguished operators whose descendants can be used to deform the holomorphic-topological theory to two theories that are fully topological. Roughly speaking, this deformation can be thought of as turning on a suitable superpotential. From a BV/BRST perspective, this deforms the differential $Q$ of the $HT$-twisted theory, which is the sum of the BV/BRST differential and the twisting supercharge $Q=Q_{BRST}+Q_{HT}$, to $Q_{A/B}=Q+\delta_{A/B}$. The effect of such a superpotential deformation on the boundary vertex algebra can be understood using the analysis of CDG . As we review below, see also twistedN=4 ; GRW , the second operator plays a distinguished role after deforming: it trivializes (makes homotopically trivial) the bulk’s dependence on the holomorphic coordinate and oftentimes realizes an action of the Virasoro algebra on boundary local operators (thus producing a VOA). Importantly, the form of the superpotential used to deform to topological theories instructs which boundary conditions are deformable in the sense of CG ; BLS .
@@
where we have $r$ copies of a $U(1)$ gauge theory with a chiral multiplet of charge $1$, a suitable matrix $K_{r}$ of mixed Chern-Simons level, and $V_{{\mathfrak{m}}_{i}}$ are a basis of gauge-invariant half-BPS monopole operators. In the IR, a $U(1)\subset(U(1)^{r})_{\text{top}}$ subgroup of the topological symmetry mixes with the $\mathcal{N}=2$ R-symmetry $U(1)_{R}$ to form the $\mathcal{N}=4$ $R$-symmetry group $SU(2)_{H}\times SU(2)_{C}$; the remaining topological symmetries are broken by the monopoles operators, in agreement with the rank-0 expectation. One of the results of GKS was that certain half indices of these theories, in the limits appropriately reproducing the $A$ and $B$ twists, appear to be vector-valued modular forms closely related to characters of the Virasoro minimal models $M(2,2r+3)$. As we now explain, our main motivation to study these boundary VOAs came from the observation of these modular phenomena as well as the appearance of Virasoro minimal models. We note that these minimal models also arise in the context of 4d Argyres-Douglas theories Beem:2017ooy and their reduction to 3d Dedushenko:2018bpp ; MD:ToAppear .
@@
    Based on examples Costello:2018swh , the Higgs branch conjecture of Beem-Rastelli in 4d, as well as free-field realizations in 3d by Beem and the first author Beem:2023dub , the Higgs/Coulomb branch chiral ring identified with (the reduced part of) Zhu’s $C_{2}$ algebra $R_{\cal V}$ of the boundary VOA ${\cal V}$; in Beem:2023dub it was conjectured that for a large class of A-twisted abelian gauge theories, the associated variety $\mathrm{Specm}(R_{\cal V})$ is indeed the Higgs branch.
@@
Thus, based on these expectations, it is natural to strengthen a bit the rank-0 conditions, and study the case where:
  1. 1)
    $R_{\cal V}$ is finite-dimensional
  2. 2)
    there are no non-trivial self-extensions of the vacuum module
The first condition, which is equivalent to saying ${\cal V}$ is $C_{2}$-cofinite, already has important consequences. In fact, it is known that under some technical assumptions (that the VOA is finitely strongly generated and non-negatively graded, which usually holds) the $C_{2}$-cofiniteness condition (and as a consequence, the zero-dimensionality of the associated variety) is equivalent to the vertex algebra being lisse arakawa2012remark , which says that the singular support of ${\cal V}$ (as a module for itself) is zero-dimensional. This further implies that its characters indeed enjoy modular properties, see e.g. Zhu . Therefore, taking seriously the first expectation means that rank-0 theories have a good chance of admitting lisse boundary VOAs, and in turn, of having half-indices that enjoy modular properties.
@@
We consider the higher-level theories ${\cal T}_{r}$ in Section 5, where a direct description of the boundary VOA becomes somewhat more difficult. We argue that there is a Dirichlet boundary condition Dir(r) that is compatible with the $B$-twist and provide evidence for the algebra of boundary local operators being identified with the simple affine VOA $L_{r}(\mathfrak{osp}(1|2))$. We finish Section 5 by comparing properties of the category of $L_{r}(\mathfrak{osp}(1|2))$ modules with the analysis of GKS , finding compatible results. The form of our half-indices can be viewed as fermionic sum representations of $L_{r}(\mathfrak{osp}(1|2))$ characters, which may be of independent interest. Finally, in Appendix A we make use of a mechanism introduced in Section 4 to propose a twisted 3d ${\cal N}=2$ TQFT and boundary condition thereof that realizes the minimal model $M(3,4)$.
@@ 3.1 The $HT$ twist of an ${\cal N}=4$ theory
We start by presenting some general features of the $HT$ twist of an ${\cal N}=4$ theory, following the conventions of twistedN=4 . The 3d ${\cal N}=4$ supersymmetry algebra has $8$ real spinor supercharges $Q^{a\dot{a}}_{\alpha}$, where $a,\dot{a}$ are Spin$(4)_{R}\simeq SU(2)_{H}\times SU(2)_{C}$ $R$-symmetry spinor indices and $\alpha$ is a $SU(2)_{E}$ Lorentz spinor index, and the following anti-commutators
@@ 3.2 Supercurrents and superpotentials
We now see how to deform the $Q_{HT}$ twist of an ${\cal N}=4$ theory to the topological $Q_{A/B}$ twist. First, we redefine the twisted spin and the $R$-charge to be those of the topological $A/B$ twist. This ensures the operator $\Theta_{A/B}$ has the necessary quantum numbers; for example, in the $A$-twist we take the $R$-charge generated by $R-S=2R_{C}$ and the twisted spin generated by $J+\tfrac{1}{2}S=R_{H}-J_{3}$. This change in twisted spin is accounted for in a modification of the higher stress tensor $G\to G_{A/B}=G\pm\tfrac{1}{2}\partial_{z}{\cal S}$. Second, we deform the action by the superpotential term:
[…]
which implements the deformation of the differential.
@@ 4 A boundary vertex algebra for ${\cal T}_{\text{min}}$
are the operators realizing the deformations to the $A$ and $B$ twist; the charge for the $U(1)_{S}$ symmetry is identified with (the negative of the) monopole number. In view of the analysis in GKS , there is a natural candidate deformable boundary condition: we expect that the boundary condition imposing
  • $(0,2)$ Dirichlet boundary conditions ${\cal D}$ on the Chern-Simons vector multiplet
  • $(0,2)$ Dirichlet boundary conditions $D$ on the chiral multiplet
is deformable to the topological $B$-twist; we call this boundary condition Dir${}=({\cal D},D)$. Indeed, Dirichlet boundary conditions for the chiral multiplet impose $\phi|=0$ and leave $\overline{\lambda}_{-}|$ unconstrained, therefore $\Theta_{B}|=0$. Moreover, the Dirichlet boundary conditions on the vector multiplet further imply the bulk local operator $\Theta_{A}$ is unconstrained on Dir. The local operators on Dir are counted by the half-index GGP ; DGP ; Yoshida:2014ssa , which is defined by
@@
Based on the discussion in Section 5, the generic Dirichlet boundary condition $({\cal D},D_{c})$ considered in GKS is similarly deformable for the $A$ twist of ${\cal T}_{\text{min}}$ – although $\Theta_{A}$ survives on Dir, for generic Dirichlet boundary conditions $D_{c}$ for chiral multiplet, which simply set $\phi|=c$, there is an additional differential that makes $\Theta_{A}$ trivial in cohomology. Moreover, the operator $\Theta_{B}$ is non-vanishing on this generic Dirichlet boundary condition, whereas it vanishes on Dir. We leave a deeper treatment of the algebra of local operators on the generic Dirichlet boundary condition of GKS for future work. [^3] Another candidate for a deformable boundary condition for the $A$ twist is $({\cal D},N)$, imposing $(0,2)$ Dirichlet boundary conditions on the vector multiplet and $(0,2)$ Neumann boundary conditions for the chiral multiplet. The Neumann boundary conditions on the chiral multiplet impose $\overline{\lambda}_{-}|=0$ which implies $\Theta_{A}=0$. However, the boundary theory has anomalous $U(1)_{H}$, which is incompatible with the $A$-twist. This can be remedied by adding a boundary fermion with $R$-charge $1$ and $U(1)_{\partial}$ charge $-1$. The $HT$-twisted half-index for $({\cal D},N)$ takes the form $I\!\!I_{({\cal D},N)}(q;y,\eta,\nu)$ $=\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{q^{\scriptsize{\frac{1}{2}}\mathfrak{m}^{2}}y^{\mathfrak{m}}[(-q^{\scriptsize{\frac{1}{2}}})^{\nu}\eta]^{\mathfrak{m}}}{(q;q)_{\infty}(yq^{\mathfrak{m}};q)_{\infty}}$ $=\bigg{(}\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{(-1)^{\mathfrak{m}}q^{\scriptsize{\frac{1}{2}}\mathfrak{m}(\mathfrak{m}+\nu)}y^{\mathfrak{m}}\eta^{\mathfrak{m}}}{(q;q)_{\infty}}\bigg{)}\bigg{(}\sum_{n\geq 0}\frac{(-1)^{n}q^{-\scriptsize{\frac{1}{2}}n(n+\nu)}\eta^{-n}}{(q;q)_{n}}\bigg{)}\ .$ Deforming to the $A$-twist amounts to sending $\eta\to 1$ and $\nu\to-1$, leading to $I\!\!I_{A}(q;y)=I\!\!I_{({\cal D},N)}(q;y,1,-1)=\bigg{(}\sum_{\mathfrak{m}\in\mathbb{Z}}\frac{(-1)^{\mathfrak{m}}q^{\scriptsize{\frac{1}{2}}\mathfrak{m}(\mathfrak{m}-1)}y^{\mathfrak{m}}}{(q;q)_{\infty}}\bigg{)}\bigg{(}\sum_{n\geq 0}\frac{(-1)^{n}q^{\scriptsize{\frac{1}{2}}n(1-n)}}{(q;q)_{n}}\bigg{)}\ ,$ which has unbounded powers of $q$. If we add a boundary Fermi multiplet with $U(1)_{\partial}$ and $U(1)_{R}$ charges $-1$ and $1$ respectively, as required to cancel the boundary anomaly, we need to multiply $\mathfrak{m}$th summand in the half-index by $(yq^{\mathfrak{m}};q)_{\infty}(q^{1-\mathfrak{m}}y^{-1};q)_{\infty}~{}.$ This has the pleasing effect of making the $q$-expansion bounded from below.
@@ 4.1.1 The $HT$-twisted boundary vertex algebra
The OPE of the abelian current $B$ with itself reflects the effective Chern-Simons $k_{eff}=2$, the OPE between the current $B$ and the fermion $\overline{\lambda}$ encodes the fact that it has gauge charge $-1$, and the OPE of the fermion $\overline{\lambda}$ with itself is regular due to the absence of a bulk superpotential, cf. Eq. (5.28) of CDG . Each of these operators has spin $1$.
@@
$\sigma_{\mathfrak{m}}(B_{n})=B_{n}-2\mathfrak{m}\delta_{n,0}\ ,\qquad\sigma_{\mathfrak{m}}(\overline{\lambda}_{n})=\overline{\lambda}_{n+\mathfrak{m}}\ ,$ (40)
@@
Denoting the operators corresponding to the states $\sigma_{\mathfrak{m}}(|0\rangle)$ by $v_{\mathfrak{m}}(z)$, these expressions translate to the following OPEs:
$B(z)v_{\mathfrak{m}}(w)\sim\frac{2\mathfrak{m}v_{\mathfrak{m}}(w)}{z-w}\qquad\overline{\lambda}(z)v_{\mathfrak{m}}(w)\sim(z-w)^{\mathfrak{m}}\sigma_{\mathfrak{m}}(\lambda_{-1}|0\rangle)(w)+\dots$ (43)
The OPEs of the boundary monopoles $v_{\mathfrak{m}}$’s with one another are simply those of ordinary vertex operators:
$v_{\mathfrak{m}_{1}}(z)v_{\mathfrak{m}_{2}}(w)=(z-w)^{2\mathfrak{m}_{1}\mathfrak{m}_{2}}:\!v_{\mathfrak{m}_{1}}(z)v_{\mathfrak{m}_{2}}(w)\!:\sim(z-w)^{2\mathfrak{m}_{1}\mathfrak{m}_{2}}v_{\mathfrak{m}_{1}+\mathfrak{m}_{2}}(w)+\dots$ (44)
We note that the subalgebra generated by $B$ and the boundary monopoles $v_{\mathfrak{m}}$ can be identified with a rank-1 lattice VOA at level $2$, i.e. based on the lattice $\sqrt{2}\mathbb{Z}$. This particularly important lattice VOA can be identified with the symmetry algebra of a compact boson at its critical radius: (the simple quotient of) an affine $\mathfrak{sl}(2)$ current algebra at level $1$. In particular, if we identify $B\leftrightarrow h$, $e\leftrightarrow V_{1}$, and $f\leftrightarrow V_{-1}$, Eq. (44) translates to the following OPEs:
@@
Equation (43) implies that that the OPE of $f$ and $\overline{\lambda}$ is non-singular, whence it is a lowest weight vector of a doublet for this affine $\mathfrak{sl}(2)$ symmetry; the state $\overline{\lambda}_{0}\sigma_{1}(|0\rangle)=\sigma_{1}(\overline{\lambda}_{-1}|0\rangle)$ fills out the remaining states in this doublet and we identify it with the restriction of the bare bulk monopole $V_{1}|$. (The bare bulk monopole $V_{-1}|$ is identified with $f$.) If we denote the operator dual to this state $\theta_{+}$ and rename $\overline{\lambda}=\theta_{-}$, Eq. (43) takes the form
@@
The remaining OPEs of the fermionic and bosonic generators ultimately follow from the affine $\mathfrak{sl}(2)$ symmetry. More explicitly, the OPEs of $\theta_{\pm}$ and $h$ follow from Eq. (38) and (43):
$h(z)\theta_{\pm}(w)\sim\frac{\pm\theta_{\pm}(w)}{z-w}$ (47)
Using the fact that $|\theta_{+}\rangle=-e_{0}|\theta_{-}\rangle$ together with $f_{n}|\theta_{-}\rangle=0$ and $h_{n}|\theta_{-}\rangle=-\delta_{n,0}|\theta_{-}\rangle$ for all $n\geq 0$, we conclude
$f_{n}|\theta_{+}\rangle=h_{n}|\theta_{-}\rangle=-\delta_{n,0}|\theta_{-}\rangle\rightsquigarrow f(z)\theta_{+}(w)\sim\frac{-\theta_{-}(w)}{z-w}$ (48)
That the OPE of $e$ and $\theta_{+}$ is non-singular follows from the fact that $e_{n}|\theta_{-}\rangle=0$ for all $n>0$ and that $e_{0}^{2}|\theta_{-}\rangle=0$ because there are no states with the necessary charge an spin. Finally, we note that the OPE of $\theta_{+}$ and $\theta_{-}$ is non-singular; this follows from the action of $\overline{\lambda}_{n}$ on $\sigma_{1}(\overline{\lambda}_{-1}|0\rangle)$. We expect that the algebra of boundary local operators is (strongly, but not freely) generated by the bosonic operators $e,f,h$ and the fermionic operators $\theta_{\pm}$.